OpenStudy (anonymous):

what annual nominal rate compounded monthly has the same annual percentage yield as 8% compounded continuously

OpenStudy (amistre64):

read over something

OpenStudy (amistre64):

e^(rt) = 1+x/12 that might be better

OpenStudy (amistre64):

e^(.08) = 1+x/12 e^(.08) - 1 = x/12 12e^(.08) - 12 = x

OpenStudy (anonymous):

i do not understand hwo to do this on my graphing calculator can you please help me

OpenStudy (dumbcow):

he forgot to raise it to 12th power since you are comparing annual rates \[(1+\frac{x}{12})^{12} = e^{.08}\] \[1+\frac{x}{12} = e^{.08/12}\] \[x = 12(e^{.08/12}-1)\]

OpenStudy (dumbcow):

your calculator has an "e" button, then input the exponent ".08/12" in parenthesis

OpenStudy (anonymous):

is thee answer .080267

OpenStudy (dumbcow):

yes

OpenStudy (anonymous):

thank you soooooooooooooo much foryour help

OpenStudy (dumbcow):

yw